OpenStudy (anonymous):

What is the slope of the line through the points (-2, 4) and (-1, -1)?

OpenStudy (anonymous):

Slope is \[=\frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

Slope is \[\frac{ -1 - (-2) }{ -1 - 4 }\]

OpenStudy (anonymous):

slope m= y2-y1/x2-x1

OpenStudy (anonymous):

Reduced slope is \[\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

ok. so my answer would be 3/5?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thankyou guys